OpenStudy (anonymous):

Approximate the real number solutions to the polynomial function f(x) = x3 - 4x2 - 4x + 16 .

OpenStudy (anonymous):

x^3 - 4x^2 - 4x + 16 x^2 (x-4) - 4 (x-4) (x^2-4) (x-4) (x+4)(x-4)(x-4)

OpenStudy (anonymous):

x = -4, x = -2, x = 2 x = -2, x = 2, x = 4 x = -4, x = -2, x = 4 x = -4, x = 2, x = 4

OpenStudy (anonymous):

oh fk i messed up x^3 - 4x^2 - 4x + 16 = 0 x^2 (x-4) - 4 (x-4) = 0 (x^2-4) (x-4) = 0 (x-2)(x+2)(x-4) = 0

OpenStudy (anonymous):

hahaha so 2,-2, 4? makes more sence haha ty